OpenStudy (anonymous):

Please Help Two people each roll a die. The probability at least one rolls a 6 is _?

OpenStudy (anonymous):

Two people each roll a die or roll a dice once is same P(6)=1/6 P(not rolling 6 on one die)=1-1/6=5/6 P(not rolling 6 on two dies )=5/6*5/6=25/36 P(at least one 6 on two dies)=1-25/36=?